# Gait Transitions in Load-Pulling Quadrupeds: Insights from Sled Dogs and a Minimal SLIP Model

### v1

@@ V Discussion and Conclusion
Among the parameters examined, swing leg stiffness emerged as a dominant control variable for enabling gait transitions. Biomechanically, this parameter governs how rapidly a leg swings through the air: higher stiffness corresponds to faster oscillation and earlier touchdown. Our results demonstrate that targeted modulation of swing stiffness in just one limb pair was sufficient to replicate the experimentally observed transitions. Specifically, modulating hindlimb stiffness alone reproduced the transition between TR and RL, while adjusting forelimb stiffness enabled the transition from RR to TR. These adjustments effectively alter the sequencing of leg touchdown events and thus the overall footfall pattern. Sensitivity analyses further confirmed that the cost function was highly responsive to perturbations in the switched limb pair and relatively insensitive to changes in the other limbs. This suggests that sled dogs may exploit stride-wise adjustments in swing timing—achieved through limb-specific stiffness control—as a localized, efficient strategy to switch gaits without global reconfiguration or speed change.
@@
Although our model replicates footfall timings and tugline dynamics with high accuracy, the underlying causes for the observed transitions remain unresolved. One hypothesis is that transitions are influenced by fluctuations in tugline force. However, across 120 RL strides from Individual 1 and 179 TR strides from Individual 2, most strides—79% and 99% respectively—did not involve a transition. While some transitions coincided with observable force changes, others occurred without any clear variation, as seen in Fig. 6 and Fig. 7. Due to the small number of transitions, statistical evaluation is not currently feasible. Larger datasets with more frequent gait changes and better-resolved force data would help clarify whether load dynamics play a causal role.
One possible explanation is that variations in pulling force influence the energetic interaction between the dog and the sled, thereby prompting gait transitions. Evaluating this hypothesis is limited by the available data. Among 120 rotary-left (RL) strides from Individual 1 and 179 transverse-right (TR) strides from Individual 2, over 79% and 99% of strides remained unchanged, respectively. While changes in tugline force were observed in some transition instances, as shown in Fig. 6 and Fig. 7, similar transitions also occurred with little to no change in force. Again, additional data with more frequent gait switches would be necessary to determine whether load dynamics are a causal factor in gait selection.
@@ Acknowledgements.
We gratefully acknowledge the generous support of the sled dog mushers who volunteered their kennel and dogs for data collection. Their willingness to share their time, expertise, and home made this fieldwork possible. Funding: S.B. acknowledges funding from NSF CAREER IOS-1941933 and Schmidt Sciences, LLC. Author contributions: Conceptualization: J.D., B.S. Experiments and related data analysis: B.S. Model and related analysis: J.D. Writing—original draft: J.D., B.S. Writing—review and editing: J.D., B.S., S.B., Z.G. Visualization: J.D., B.S. Supervision, funding acquisition: S.B., Z.G. Competing interests: The authors declare no competing interests. Data and materials availability: All code to run simulations, generate figures, and reproduce results is publicly available at ….

### v2

@@ V Discussion and Conclusion
Among the parameters examined, swing leg stiffness emerged as a dominant control variable for enabling gait transitions. Biomechanically, this parameter governs how rapidly a leg swings through the air: higher stiffness corresponds to faster oscillation and earlier touchdown. Our results demonstrate that targeted modulation of swing stiffness in just one limb pair was sufficient to replicate the experimentally observed transitions. Specifically, modulating hindlimb stiffness alone reproduced the transition between TranR and RotL, while adjusting forelimb stiffness enabled the transition from RotR to TranR. These adjustments effectively alter the sequencing of leg touchdown events and thus the overall footfall pattern. Sensitivity analyses further confirmed that the cost function was highly responsive to perturbations in the switched limb pair and relatively insensitive to changes in the other limbs. This suggests that sled dogs may exploit stride-wise adjustments in swing timing—achieved through limb-specific stiffness control—as a localized, efficient strategy to switch gaits without global reconfiguration or speed change.
@@
Although our model replicates footfall timings and tugline dynamics with high accuracy, the underlying causes for the observed transitions remain unresolved. One hypothesis is that transitions are influenced by fluctuations in tugline force. However, across 120 RotL strides from Individual 1 and 179 TranR strides from Individual 2, most strides—79% and 99% respectively—did not involve a transition. While some transitions coincided with observable force changes, others occurred without any clear variation, as seen in Fig. 6 and Fig. 7. Due to the small number of transitions, statistical evaluation is not currently feasible. Larger datasets with more frequent gait changes and better-resolved force data would help clarify whether load dynamics play a causal role.
One possible explanation is that variations in pulling force influence the energetic interaction between the dog and the sled, thereby prompting gait transitions. Evaluating this hypothesis is limited by the available data. Among 120 rotary-left (RotL) strides from Individual 1 and 179 transverse-right (TranR) strides from Individual 2, over 79% and 99% of strides remained unchanged, respectively. While changes in tugline force were observed in some transition instances, as shown in Fig. 6 and Fig. 7, similar transitions also occurred with little to no change in force. Again, additional data with more frequent gait switches would be necessary to determine whether load dynamics are a causal factor in gait selection.
@@ Acknowledgements.
We gratefully acknowledge the generous support of the sled dog mushers who volunteered their kennel and dogs for data collection. Their willingness to share their time, expertise, and home made this fieldwork possible. Funding: S.B. acknowledges funding from NSF CAREER IOS-1941933 and Schmidt Sciences, LLC. Author contributions: Conceptualization: J.D., B.S. Experiments and related data analysis: B.S. Model and related analysis: J.D. Writing—original draft: J.D., B.S. Writing—review and editing: J.D., B.S., S.B., Z.G. Visualization: J.D., B.S. Supervision, funding acquisition: S.B., Z.G. Competing interests: The authors declare no competing interests. Data and materials availability: All code to run simulations, generate figures, and reproduce results is publicly available at our online repository .